OpenStudy (anonymous):

Given a polynomial f(x), if (x - 1) is a factor, what else must be true?

OpenStudy (science0229):

If (x-1) is a factor of f(x), f(x) can be written as (x-1)*g(x) where g(x) is another polynomial.

OpenStudy (science0229):

If x=1, you get f(1)=(1-1)*g(1)=0*g(1)=0 So, you get that f(1)=0 is true.

OpenStudy (science0229):

Just for bonus, if f(x) has (x-c) as its factor (c=constant), f(c)=0 is always true.

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

Better known as the Factor Theorem.